# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ III.3 Deriving the sensitivity and thermal power
The derivations are performed for the main transition under consideration, i.e., the $27^{2}\text{D}_{5/2}\rightarrow 26^{2}\text{F}_{7/2}$ transition at $125.836\text{\,}$. Photon counts with acquisition times of $1\text{\,}$ are measured for $30\text{\,}$ and averaged to get the mean counts of $(9.87\pm 0.25)\text{\times}{10}^{3}\text{\,}\mathrm{p}$hotons per second. Noticeably, the statistics are super-Poissonian, which is expected for thermal light, but for single shot measurements of multimode light at a measurement time far exceeding the bandwidth inverse, the quantum phenomena of thermal bunching should cancel out. What is instead observed are the macroclassical gain fluctuations of the setup, which can be calculated to constitute a relative error of about $1.5\text{\,}$. Shorter measurements are more likely to approach the shot noise (Poissonian) limit, as well as mitigate these macro effects, however, this is not a priority of this work.
Following the calibration stage, the converted photon counts can be translated into the incident electric field of $E_{\text{n}}=$36.7\pm 0.7\text{\,}$\$ and further into thermal field intensity through $I=E_{\text{n}}^{2}/Z_{0}=$3.57\pm 0.13\text{\,}$\$, where $Z_{0}\approx$377\text{\,}$\$ is the impedance of free space. Multiplying this result by the small antenna aperture $A$ defined as:
@@
with the antenna gain $g=9$ obtained from field phase-matching considerations, the isotropic power of the thermal radiation equal to $145\pm 5.\text{\,}$ can be retrieved.
Finally, the parameters of interest can be calculated. Sensitivity is derived as the thermal intensity scaled by the ratio of thermal photon count fluctuations to the mean counts and calculated as $90.3\pm 3.4\text{\,}$. Consequently, the NEP equals to sensitivity multiplied by $A$, yielding a value of $3.67\pm 0.14\text{\,}$.
@@ III.4 Deriving the frequency comb mode number
Observing particular teeth and performing frequency scans happens through detuning from the resonant case, which happens in relative units. By estimating the mode number, we can move to absolute frequencies, due to the information embedded in the comb peak separation. Knowing that the mode number is an integer, it is possible to infer the transition frequency, which for some transitions can be further verified using the radar chip. As it is stabilized using a phase-locked loop, it has a known frequency and can confirm the validity of the derivations.
We have determined the MaiTai repetition rate $f_{\text{rep}}$ to be $79.490\,20\pm 0.000\,01\text{\,}$ using Agilent N9010A spectrum analyzer. In the case of the main transition under consideration, an ARC value of $125.836\pm 0.004\text{\,}$ is calculated, with the uncertainty obtained by redoing the calculations and propagating the uncertainty of quantum defects [41, 42]. The next step involves locating the working point, at which the scan visible in Fig. 3 a) is performed. The thermal profile is equivalent to integral over the THz frequencies, so by fitting a Lorentzian profile the resonant $\Delta_{1200}$ value can be found. This particular model shows excellent agreement with the measured thermal photon data, indicating that the resonance is at $\Delta_{1200}=$4.92\pm 0.05\text{\,}$\$.
@@
Given the small, but still relevant deviation from the value provided by ARC, an additional verification is performed using the radar chip source. As it is controlled by a phase-locked loop with a known reference frequency, we are able to assess the frequency of the comb tooth to be located at a target frequency of $125.829\pm 0.001\text{\,}$ corresponding to a mode number estimate of $1582.95\pm 0.04\text{\,}$, confirming the previous result. The parameters, used in the derivation of tooth mode number, are presented in Tab. 1.
@@ III.5 Beating of adjacent frequency comb peaks
Further evidence of frequency comb detection can be presented by measuring the beat notes between individual peaks. As they are separated by multiples of $f_{\text{rep}}$, the beat note frequencies will also assume values of $nf_{\text{rep}}$. The beating manifests itself in the time domain, so we observe the time resolved photon conversion response on the timetagger, connected to the single photon counter. The 40 MHz MaiTai internal photodiode readout is used as the external trigger of the timetagger, allowing for measurement of all the potential beat-note frequencies.
As with the frequency scans, detuning the decoupling laser allows adjusting the sensitivity to different frequencies and a value that sits between two comb peaks is selected so as to make the beatings more pronounced in comparison to the resonant case.
@@
Figure 5: a) Temporal profile of the measured conversion signal, corresponding to mean-normalized photon counts. The blue lines correspond to the frequency comb resonant peak condition, while the case for red lines is detuned by $\Delta_{1200}=$40\text{\,}$\$. Dashed lines represent results obtained from applying Eq. (6) to the fitted conversion gain. b)-c) The Fourier transforms of the respective temporal profiles, with the resonant case having significantly lower between oscillating terms and constant contribution. Again, dashes represent theoretical predictions, with the signal amplitudes scaled to experimental data levels.
@@
where $A_{n}$ is the amplitude of the $n$-th frequency component, and $\omega_{n}=2\pi nf_{\text{rep}}$. For the convenience of notation, we denote $\mathcal{G}(\Delta_{\text{THz}}+nf_{\text{rep}},\Delta_{1200})$ as $\mathcal{G}_{n}$. The equation above can be split into two parts, namely the non-oscillating part, described by Eq. (2), and the time-dependent part given as:
@@ IV Discussion
The obtained THz intensity detection NEP values show an improvement of 2 to 4 orders of magnitude over state of the art room temperature alternatives [43, 44, 45, 46, 47, 48, 49], showcasing great potential for Rydberg atoms as an alternative platform in this field. Similar conclusions can be drawn from other attempts at using a Rydberg atom based converter for THz detection [50], where, we also achieved an improvement of about 4 orders of magnitude in sensitivity. Direct comparison of NEP is difficult in this case, due to different assumptions in aperture derivation.
@@ Acknowledgments
We thank M. Lipka for his initial contribution to the construction of the pulsed THz source.
This research was funded in whole or in part by National Science Centre, Poland grants No. 2021/43/D/ST2/03114 and No. 2024/53/N/ST7/02730. The "Quantum Optical Technologies" (FENG.02.01-IP.05-0017/23) project is carried out within the Measure 2.1 International Research Agendas programme of the Foundation for Polish Science co-financed by the European Union under the European Funds for Smart Economy 2021-2027 (FENG).
@@
To determine the attenuation imposed by the HWP combined with the polarizer, the transmission of such pair needs to be calibrated as a function of the HWP rotation angle, for which we use the photon counts measurement as a function of angles of both HWP for the $I_{2}$ TRA source intensity. For the sake of calibration of each of the individual waveplates, we choose the cross-sections of measured photon counts through the rotation angle of the other waveplate corresponding to its measured maximum transmission. As the HWPs are not fully rigid and could move by about $\pm 0.5^{\circ}$ on their own, the uncertainty of the transmission functions is taken as a difference between the interpolated function and the interpolated function shifted by $0.5^{\circ}$. The interpolated transmission functions together with the data points and uncertainties can be seen in Fig. 6. Additionally, based on found transmission functions, maximum attenuations of both pairs can be estimated as $A_{1}=$19.6\pm 1.7\text{\,}$\$ and $A_{2}=$23.6\pm 0.4\text{\,}$\$ for the first and second pair respectively.
@@
where $H$ is the Hamiltonian of the system, $\rho$ is the density matrix and $\mathcal{L}_{\Gamma}$ is a Lindblad superoperator, which accounts for decoherence mechanisms. We model five-level atomic ladder in the RWA (rotating-wave approximation) ,which produces the following form of the Hamiltonian in dipolar approximation:
$H=-\tfrac{\hbar}{2}\begin{pmatrix}\scriptstyle 0&\scriptstyle\Omega^{*}_{780}&0&0&0\\ \scriptstyle\Omega_{780}&\scriptstyle i\Gamma_{1}+2\Delta_{\textbf{I}}&\scriptstyle\Omega^{*}_{480}&0&0\\ 0&\scriptstyle\Omega_{480}&\scriptstyle i\Gamma_{2}+2\Delta_{\textbf{II}}&\scriptstyle\Omega^{*}_{\text{THz}}&0\\ 0&0&\scriptstyle\Omega_{\text{THz}}&\scriptstyle i\Gamma_{3}+2\Delta_{\textbf{III}}&\scriptstyle\Omega^{*}_{1200}\\ 0&0&0&\scriptstyle\Omega_{1200}&\scriptstyle i\Gamma_{4}+2\Delta_{\textbf{IV}}\end{pmatrix}-\tfrac{i\Gamma_{\text{tr}}}{2}\mathbb{I},$ (8)
@@
Room temperature atomic vapors are significantly influenced by the Doppler effect, stemming from the atoms moving along the electric field propagation axis. Factoring in the broadening of the resonance can be simply implemented as additional diagonal values of the Hamiltonian. For that, we define the $K$ matrix, whose diagonal elements represent cumulative sums of the wave vectors of the fields required to excite the system to the corresponding energy levels, with each subsequent level involving an additional driving field. The signs are picked to reflect the counter propagation of the probe beam, yielding:
@@
and obtain the final solution by averaging over the Maxwell-Boltzmann velocity distribution, represented by a statistic weight of $w(v)$:
@@
IV.1 Modeling of Autler-Townes splitting
@@
IV.2 Modeling of terahertz frequency comb
@@
As mentioned in the main text, for the sake of the calibration of the receiver, the Rabi frequency of the THz field is determined based on the measured A-T splitting visible in the probe transmission spectra. The values of the Rabi frequencies are found by numerically fitting the steady-state solution for the theoretical model described in Appendix B to the probe transmission signal. In such fit, the Rabi frequency of the THz field is treated as one of the parameters of the fitted function. Exemplary probe transmission spectra, together with the fitted theoretical model, can be seen in Fig. 7. In order to perform the fitting properly, both the experimental and numerical data are min-max normalized. To further increase the accuracy of the numerical fits, the common parameters, which remained constant during the measurement, such as the Rabi frequencies of laser fields, are fitted simultaneously to all the measured spectra.
@@
The calculated conversion gain can be seen in Fig. 8, showing two resonances that give a good working intuition for performing comb spectroscopy with this setup. A single photon count measurement, given a frequency comb and a fixed value of $\Delta_{1200}$, can be understood as a discrete sum of conversion gain values along a vertical line of the map, as expressed in Eq. (2). The frequency comb teeth remain static when tuning the decoupling laser, meaning that we adjust the conversion gain to the positions of the teeth. The resonances observed in Fig. 3 are spaced at $f_{\text{rep}}$, because we count two teeth at this interval. On the other hand, the thermal photon measurements, correspond to taking an integral across the $y$-axis.
Figure 8: Map of $|\mathcal{G}|^{2}$ with respect to $\Delta_{\text{THz}}$ and $\Delta_{1200}$. It is directly proportional to the conversion signal intensity, and as such the color map is trimmed at the order of magnitude corresponding to the noise floor. The comb teeth are marked on the vertical axis and can be thought of as sampling points for a given value of $\Delta_{1200}$.
@@
The sensor’s parameters, used in the considerations of thermal noise level, as well as sensor’s sensitivity an NEP, are presented in Tab. 2.
@@
Thermal intensity $3.57\pm 0.13\text{\,}$
Thermal power $145\pm 5.\text{\,}$
Sensitivity $90.3\pm 3.4\text{\,}$
NEP $3.67\pm 0.14\text{\,}$

### v2

@@ III.3 Deriving the sensitivity and thermal power
[…]
Following the calibration stage, presented in Fig. 2, the converted photon counts can be translated into the equivalent incident electric field of $E_{\text{noise}}=$36.7\pm 0.7\text{\,}$\$ and further into intensity through $I_{\text{noise}}=E_{\text{noise}}^{2}/2Z_{0}=$1.78\pm 0.07\text{\,}$\$, where $Z_{0}\approx$377\text{\,}\text{Ohm}$\$ is the impedance of free space. Next, to obtain NEP, the receiver’s effective aperture needs to be considered. For this, we can consider the gain-defined small antenna aperture $A$ as:
@@
where $g$ is the antenna gain (directivity) and $\lambda=$2.38\text{\,}$\$ is the wavelength of received field. We obtain the antenna gain $g=9$ from field phase-matching considerations (i.e., the receiver reception pattern) [36], which leads to the result of $A=$4.07\text{\,}\mathrm{m}\mathrm{m}^{2}$\$. Consequently, the effective registered power corresponding to the thermal radiation equal to $P_{\text{noise}}=I_{\text{noise}}A=$72.5\pm 2.7\text{\,}$\$ can be retrieved.
[…]
@@ III.4 Deriving the frequency comb mode number
Observing particular teeth and performing frequency scans happens through detuning from the resonant case, which happens in relative units. By estimating the mode number, we can move to absolute frequencies, due to the information embedded in the comb peak separation. Knowing that the mode number is an integer, it is possible to infer the transition frequency, which, for some transitions, can be further verified using the radar chip. As it is stabilized using a phase-locked loop, it has a known frequency and can confirm the validity of the derivations.
We have determined the MaiTai repetition rate $f_{\text{rep}}$ to be $79.490\,20\pm 0.000\,01\text{\,}$ using Agilent N9010A spectrum analyzer. In the case of the main transition under consideration, an ARC value of $125.836\pm 0.004\text{\,}$ is calculated, with the uncertainty obtained by redoing the calculations and propagating the uncertainty of quantum defects [41, 42]. The next step involves locating the working point, at which the scan visible in Fig. 3 a) is performed. The thermal profile is equivalent to an integral over the THz frequencies, so by fitting a Lorentzian profile, the resonant $\Delta_{1200}$ value can be found. This particular model shows excellent agreement with the measured thermal photon data, indicating that the resonance is at $\Delta_{1200}=$4.92\pm 0.05\text{\,}$\$.
@@
Given the small but still relevant deviation from the value provided by ARC, an additional verification is performed using the radar chip source. As it is controlled by a phase-locked loop with a known reference frequency, we are able to assess the frequency of the comb tooth to be located at a target frequency of $125.829\pm 0.001\text{\,}$ corresponding to a mode number estimate of $1582.95\pm 0.04\text{\,}$, confirming the previous result. The parameters used in the derivation of tooth mode number are presented in Tab. 1.
@@ III.5 Beating of adjacent frequency comb peaks
Further evidence of frequency comb detection can be presented by measuring the beat notes between individual peaks. As they are separated by multiples of $f_{\text{rep}}$, the beat note frequencies will also assume values of $nf_{\text{rep}}$. The beating manifests itself in the time domain, so we observe the time-resolved photon conversion response on the timetagger, connected to the single photon counter. The 40 MHz MaiTai internal photodiode readout is used as the external trigger of the timetagger, allowing for measurement of all the potential beat-note frequencies.
As with the frequency scans, detuning the decoupling laser allows adjusting the sensitivity to different frequencies and a value that sits between two comb peaks are selected so as to make the beatings more pronounced in comparison to the resonant case.
@@
Figure 5: a) Temporal profile of the measured conversion signal, corresponding to mean-normalized photon counts. The blue lines correspond to the frequency comb resonant peak condition, while the case for red lines is detuned by $\Delta_{1200}=$40\text{\,}$\$. Dashed lines represent results obtained from applying Eq. (6) to the fitted conversion gain. b)-c) The Fourier transforms of the respective temporal profiles, with the resonant case having significantly lower contrast between the oscillating terms and the constant contribution. Again, dashes represent theoretical predictions, with the signal amplitudes scaled to experimental data levels. Contrary to a), the full signal is presented here.
@@
where $A_{n}$ is the amplitude of the $n$-th frequency component, and $\omega_{n}=2\pi nf_{\text{rep}}$. For the convenience of notation, we denote the conversion gain $\mathcal{G}(\Delta_{\text{THz}}+nf_{\text{rep}},\Delta_{1200})$ as $\mathcal{G}_{n}$. The equation above can be split into two parts, namely the non-oscillating part, described by Eq. (2), and the time-dependent part given as:
@@ IV Discussion
The obtained THz intensity detection NEP values show an improvement of 2 to 4 orders of magnitude over state of the art room temperature alternatives [43, 44, 45, 46, 47, 48, 49], showcasing great potential for Rydberg atoms as an alternative platform in this field. Similar conclusions can be drawn from other attempts at using a Rydberg atom-based converter for THz detection [50], where we also achieved an improvement of about 4 orders of magnitude in sensitivity. Direct comparison of NEP is difficult in this case, due to different assumptions in aperture derivation.
@@ Acknowledgments
We thank M. Lipka for his initial contribution to the construction of the pulsed THz source. This research was funded in whole or in part by the National Science Centre, Poland, grants No. 2021/43/D/ST2/03114 and No. 2024/53/N/ST7/02730. The "Quantum Optical Technologies" (FENG.02.01-IP.05-0017/23) project is carried out within the Measure 2.1 International Research Agendas programme of the Foundation for Polish Science, co-financed by the European Union under the European Funds for Smart Economy 2021-2027 (FENG). S.B. is a recipient of the Foundation for Polish Science START 2025 scholarship.
@@
To determine the attenuation imposed by the HWP combined with the polarizer, the transmission of such a pair needs to be calibrated as a function of the HWP rotation angle, for which we use the photon counts measurement as a function of angles of both HWP for the $I_{2}$ TRA source intensity. For the sake of calibration of each of the individual waveplates, we choose the cross-sections of measured photon counts through the rotation angle of the other waveplate corresponding to its measured maximum transmission. As the HWPs are not fully rigid and could move by about $\pm 0.5^{\circ}$ on their own, the uncertainty of the transmission functions is taken as a difference between the interpolated function and the interpolated function shifted by $0.5^{\circ}$. The interpolated transmission functions together with the data points and uncertainties can be seen in Fig. 6. Additionally, based on the found transmission functions, maximum attenuations of both pairs can be estimated as $A_{1}=$19.6\pm 1.7\text{\,}$\$ and $A_{2}=$23.6\pm 0.4\text{\,}$\$ for the first and second pair, respectively.
@@
where $H$ is the Hamiltonian of the system, $\rho$ is the density matrix, and $\mathcal{L}_{\Gamma}$ is a Lindblad superoperator, which accounts for decoherence mechanisms. We model a five-level atomic ladder in the RWA (rotating-wave approximation), which produces the following form of the Hamiltonian in dipolar approximation:
$H=-\tfrac{\hbar}{2}\begin{pmatrix}\scriptstyle 0&\scriptstyle\Omega^{*}_{780}&0&0&0\\ \scriptstyle\Omega_{780}&\scriptstyle i\Gamma_{1}+2\Delta_{\textbf{I}}&\scriptstyle\Omega^{*}_{480}&0&0\\ 0&\scriptstyle\Omega_{480}&\scriptstyle i\Gamma_{2}+2\Delta_{\textbf{II}}&\scriptstyle\Omega^{*}_{\text{THz}}&0\\ 0&0&\scriptstyle\Omega_{\text{THz}}&\scriptstyle i\Gamma_{3}+2\Delta_{\textbf{III}}&\scriptstyle\Omega^{*}_{1200}\\ 0&0&0&\scriptstyle\Omega_{1200}&\scriptstyle i\Gamma_{4}+2\Delta_{\textbf{IV}}\end{pmatrix}-\tfrac{i\hbar\Gamma_{\text{tr}}}{2}\mathbb{I},$ (8)
@@
Room temperature atomic vapors are significantly influenced by the Doppler effect, stemming from the atoms moving along the electric field propagation axis. Factoring in the broadening of the resonance can be simply implemented as additional diagonal values of the Hamiltonian. For that, we define the $K$ matrix, whose diagonal elements represent cumulative sums of the wave vectors $k$ of the fields required to excite the system to the corresponding energy levels, with each subsequent level involving an additional driving field. The signs are picked to reflect the counter propagation of the probe beam, yielding:
@@
where $v$ denotes the velocity class of atoms, and obtain the final solution by averaging over the Maxwell-Boltzmann velocity distribution, represented by a statistical weight of $w(v)$:
@@
Modeling of Autler-Townes splitting
@@
Modelling of terahertz frequency comb
@@
As mentioned in the main text, for the sake of the calibration of the receiver, the Rabi frequency of the THz field is determined based on the measured A-T splitting visible in the probe transmission spectra. The values of the Rabi frequencies are found by numerically fitting the steady-state solution for the theoretical model described in Appendix B to the probe transmission signal. In such a fit, the Rabi frequency of the THz field is treated as one of the parameters of the fitted function. Exemplary probe transmission spectra, together with the fitted theoretical model, can be seen in Fig. 7. In order to perform the fitting properly, both the experimental and numerical data are min-max normalized. To further increase the accuracy of the numerical fits, the common parameters, which remained constant during the measurement, such as the Rabi frequencies of laser fields, are fitted simultaneously to all the measured spectra.
@@
The calculated conversion gain can be seen in Fig. 8, showing two resonances that give a good working intuition for performing comb spectroscopy with this setup. A single photon count measurement, given a frequency comb and a fixed value of $\Delta_{1200}$, can be understood as a discrete sum of conversion gain values along a vertical line of the map, as expressed in Eq. (2). The frequency comb teeth remain static when tuning the decoupling laser, meaning that we adjust the conversion gain to the positions of the teeth. The resonances observed in Fig. 3 are spaced at $f_{\text{rep}}$, because we count two teeth at this interval. On the other hand, the thermal photon measurements correspond to taking an integral across the $y$-axis.
Figure 8: Map of $|\mathcal{G}|^{2}$ with respect to $\Delta_{\text{THz}}$ and $\Delta_{1200}$. It is directly proportional to the conversion signal intensity, and as such, the color map is trimmed at the order of magnitude corresponding to the noise floor. The comb teeth are marked on the vertical axis and can be thought of as sampling points for a given value of $\Delta_{1200}$.
@@
The sensor’s parameters, used in the considerations of thermal noise level, as well as the sensor’s sensitivity and NEP, are presented in Tab. 2.
@@
Thermal intensity $1.78\pm 0.07\text{\,}$
Thermal power $72.5\pm 2.7\text{\,}$
Sensitivity $45.2\pm 1.7\text{\,}$
NEP $1.84\pm 0.07\text{\,}$